A sample of 5.0 mol CO2(g) is originally confined
Answer #1

ym=0.5 I= 280 K イ T. 2- I S& 3 了 γ-1 1-31-I 3

> Hi, how we learn the value of gamma ı think 1.33???

maram hallak Sat, Oct 30, 2021 8:04 AM
Answer #2
For an adiabatic process q=0 Work =-Pex ×∆V 78.5×10^³×((60-15)×0.001)) =-3532.5J or -3.5325kJ ∆T=w/n(Cp-R)=-3532.5÷5(37.11-8.315) =-24.53K ∆U=q+w but q=0 ∆U=w =-3532.5J ∆H=∆U+nR∆T =-3532.5+(5×8.314×-24.53) =-4552.40J or -4.552kJ
source: Student
answered by: Mongare